OpenStudy (anonymous):

Henry rolls 2 number cubes numbered 1 through 6 while playing his favorite board game. He will get a second turn if he rolls a sum that is an even number less than 10. What are Henry's chances of getting a second turn when he rolls the number cubes? @ganeshie8

ganeshie8 (ganeshie8):

start by calculating how many ways Henry can get an 'even sum'

OpenStudy (anonymous):

2,4,6,8,

ganeshie8 (ganeshie8):

yes, when he adds the numbers on both cubes, he should get one of them

ganeshie8 (ganeshie8):

actually we done even need to do all that calculation. just see that, there are exactly 3 even numbers and 3 odd numbers between 1 and 6

ganeshie8 (ganeshie8):

so, the chances of getting an even sum is exactly 1/2

ganeshie8 (ganeshie8):

see if that makes more or less sense

OpenStudy (anonymous):

7/18 @ganeshie8

ganeshie8 (ganeshie8):

how did u get 7/18 ?

OpenStudy (anonymous):

i dont know either lol how do we figure this out?

ganeshie8 (ganeshie8):

if you do the actual calculation you will get 18/36 = 1/2

OpenStudy (anonymous):

are you sure its not 17/36

ganeshie8 (ganeshie8):

let me think a bit

ganeshie8 (ganeshie8):

oh we missed this constraint :- less than 10.

OpenStudy (anonymous):

5/36

ganeshie8 (ganeshie8):

lets do this properly :- if he gets \(1\) on first die, then he can get \(1, 3, 5\) on second die : 3 ways if he gets \(3\) on first die, then he can get \(1, 3, 5\) on second die : 3 ways if he gets \(5\) on first die, then he can get \(1, 3, 5\) on second die : 3 ways if he gets \(2\) on first die, then he can get \(2, 4, 6\) on second die : 3 ways if he gets \(4\) on first die, then he can get \(2, 4\) on second die : 2 ways if he gets \(6\) on first die, then he can get \(2\) on second die : 1 ways add them all up

ganeshie8 (ganeshie8):

total ways it can be an even sum = 3+3+3+3+2+1 = 15

OpenStudy (anonymous):

15

ganeshie8 (ganeshie8):

so the chances are 15/36 = 5/12

ganeshie8 (ganeshie8):

u have that in ur options ha ?

OpenStudy (anonymous):

These are the options 7/18 11/18 5/36 17/36

ganeshie8 (ganeshie8):

if he gets 1 on first die, then he can get 1,3,5 on second die : 3 ways if he gets 3 on first die, then he can get 1,3,5 on second die : 3 ways if he gets 5 on first die, then he can get 1,3 on second die : 2 ways if he gets 2 on first die, then he can get 2,4,6 on second die : 3 ways if he gets 4 on first die, then he can get 2,4 on second die : 2 ways if he gets 6 on first die, then he can get 2 on second die : 1 ways Now add themm up. and do double check if we counted all the possible ways ok

OpenStudy (anonymous):

14

ganeshie8 (ganeshie8):

yes 14 is the # of ways he can get an even sum divide it by total number ways he can roll the die : 36

OpenStudy (anonymous):

7/18

OpenStudy (anonymous):

Thank You!

ganeshie8 (ganeshie8):

thats right ! do double check if we missed counting any... seems my brain is dead today lol :o

OpenStudy (anonymous):

no its not its bright as usual.

ganeshie8 (ganeshie8):

lol no, ur brain is bright, u caught my mistakes xD